OpenStudy (anonymous):

For y = -sin(x – (π / 4)) + 2, find: a. Amplitude:1 b. Period: c. Phase Shift: d. Vertical Translation:

OpenStudy (jdoe0001):

got any of those yet?

OpenStudy (anonymous):

What?

OpenStudy (jdoe0001):

hmmm ohh ahedmm I see the amplitude... is 1 ok

OpenStudy (jdoe0001):

\(\bf y=-sin\left({\color{red}{ 1}}x-{\color{blue}{ \frac{\pi}{4}}}\right){\color{brown}{ +2}} \\ \quad \\ \textit{vertical shift}\to {\color{brown}{ \square }}\qquad period\to \cfrac{2\pi}{{\color{red}{ \square }}} \\ \quad \\ \textit{phase shift}\to {\color{blue}{ \square }}\)